OpenStudy (anonymous):

w^12/w^18 = a. w^30 b. w^-30 c. w^6 d. w^-6

OpenStudy (anonymous):

Choice (D). subtract the bottom exponent (18) from the top exponent (12), so w^12/w^18 = w^ (12-18) = w^(-6)